OpenStudy (agent47):

Hey @Zarkon, sorry to bother you, have a quick question about yesterday's problem

OpenStudy (agent47):

How did you figure out that it was (x^2-x+42)/12?

OpenStudy (agent47):

I got the individual expectations and they worked correctly, but I got my answer in terms of a sum over x

OpenStudy (dan815):

whats the question pls repost

OpenStudy (agent47):

I got expected value:\[E(Y|X=x)=\frac{1}{6}(x^2+\sum_{i=x+1}^{6}{i})\]@Zarkon simplified that to \[\frac{x^2-x+42}{12}\]Just wondering how.

OpenStudy (agent47):

For x = 1, 2, 3, 4, 5, 6, I get the same answer from both, but I was just wondering how he came up with that better looking result

OpenStudy (agent47):

Teach us @Zarkon !

OpenStudy (zarkon):

\[\sum_{i=a}^{n}{i}=\sum_{i=1}^{n}{i}-\sum_{i=1}^{a-1}{i}\] \[\sum_{i=1}^{n}{i}=\frac{n(n+1)}{2}\]

OpenStudy (agent0smith):

You can shift the index to simplify it. Let n = i - x, so i = x+n Then plug i=x+1 and i=6 into n = i - x to get \[\large\sum_{i=x+1}^{6}{i} = \sum_{n=1}^{6-x}{(n+x)} \] Then you can use the summation formulas like the one zarkon gave. Remember that x is just a constant, so its sum is pretty straightforward.

OpenStudy (agent0smith):

But it's easier to shift the index if you just think of it as subtracting x from the top and bottom limits of the sum, and then adding x inside the summation \[\large\sum_{i=x+1}^{6}{i} =\sum_{i=1}^{6-x}{(i+x)} =\]Then you can break up the sum like so\[\large \sum_{i=1}^{6-x}{(i+x)} =\sum_{i=1}^{6-x}{i}+\sum_{i=1}^{6-x}{x}\] Then just use summation formulas and remember that x is a constant.

OpenStudy (agent0smith):

Just use number 1 and 2 from here http://wonderfulworldofcalulus.weebly.com/uploads/2/8/2/3/28239103/1398375903.jpg

OpenStudy (zarkon):

\[\frac{1}{6}(x^2+\sum_{i=x+1}^{6}{i})\] \[=\frac{1}{6}(x^2+\sum_{i=1}^{6}{i}-\sum_{i=1}^{x}{i})\] \[=\frac{1}{6}(x^2+\frac{6\cdot7}{2}-\frac{x(x+1)}{2})\] \[=\frac{1}{12}(2x^2+42-x(x+1))\] \[=\frac{1}{12}(x^2-x+42)\]